OpenStudy (anonymous):

Where does the graph of x2 - x - 2 = y cross the x-axis?

OpenStudy (zzr0ck3r):

when y = 0 x^2 - x - 2 = 0 x^2 +x-2x-2=0 x(x+1)-2(x+1) = 0 (x+1)(x-2)=0 x = ? and x = ?

OpenStudy (anonymous):

1 & -2?

OpenStudy (zzr0ck3r):

when x = 1 (x+1)(x-2) = (1+1)(1-2) = 2(-1) = -2

OpenStudy (zzr0ck3r):

almost break it up x+1 = 0 or x-2 = 0

OpenStudy (anonymous):

-1&2

OpenStudy (zzr0ck3r):

notice the only time two things multiplied by each other can = 0 is if one or the other is 0

OpenStudy (anonymous):

ok thanks!